OpenStudy (anonymous):

What is the x and y intercept for f(x) = - x^3 +4x

Nnesha (nnesha):

`x-intercept` is a point when line intersect the x-axis when y =0 so substitute y for 0 to find x-intercept `y-intercept` is a point when line intersect y-axis when x=0 substitute x for 0 to find y-intercept

OpenStudy (anonymous):

@ Nnesha what exactly do you mean -y^3+4y and vise versa

Nnesha (nnesha):

 to find x-intercept u should substitute y for 0 and to find y-intercept sub x for 0

OpenStudy (anonymous):

oh okay that clears it Mrs Nnesha Thank you

Nnesha (nnesha):

\[\large\rm f(x)=-x^3+4x ---- >y =-x^3 +4x\] btw f(x) is same as y

Nnesha (nnesha):

glad to help let me know if you have a question :=))

OpenStudy (anonymous):

Okie dokay